# BCR-DRL: Behavior- and Context-aware Reward for Deep Reinforcement Learning in Human-AI Coordination

### v4

@@ 4 The Proposed BCR-DRL Algorithm
The key contribution of our BCR-DRL lies in the design of our behavior- and context-aware reward (BCR). In this section, we first present the design specifics of our BCR, followed by the details of the training algorithm of BCR-DRL for HAIC.
@@ 4.1 Design Specifics of BCR
The dual intrinsic rewards encourage the exploration of critical rare state-action pairs (see Section 2) that are associated with sparse rewards, capturing distinctive behavioral patterns from both the AI agent and its human coordinator. The context-aware weights adaptively adjust each reward component’s significance epoch by epoch, maintaining synchronization with the BCR-DRL policy updates. We note that the use of distinguishable superscripts ($\mathrm{E}$ and $\mathcal{E}$) for the extrinsic reward and its context-aware weight, is intentional to emphasize their distinct design rationale, which will be elaborated in eq. (8).
@@ 4.1.1 Extrinsic Reward
The extrinsic reward, $r_{t}^{\mathrm{E}}$, comprises two components: the target sparse reward, $r_{t}^{\mathrm{E}_{\mathrm{S}}}$, and a stage reward [^1] Stage rewards are defined as the intermediate rewards used to guide the exploration and exploitation of key preliminary actions that lead to the target sparse rewards Hu et al. (2020); Ng et al. (1999). We follow the methods of stage rewards designing given in the supplementary of Carroll et al. (2019)., $r_{t}^{\mathrm{E}_{\mathrm{G}}}$. This extrinsic reward is given by
$\begin{split}r_{t}^{\mathrm{E}}&=r_{t}^{\mathrm{E}_{\mathrm{S}}}+r_{t}^{\mathrm{E}_{\mathrm{G}}}\cdot f_{\phi}(t)\\ &=\lambda^{\mathrm{E}_{\mathrm{S}}}\cdot\mathds{1}\left\{f_{C}(s_{t},a_{t})={True}\right\}+r_{t}^{\mathrm{E}_{\mathrm{G}}}\cdot f_{\phi}(t),\end{split}$ (2)
where $\lambda^{\mathrm{E}_{\mathrm{S}}}$ is a constant coefficient representing the magnitude of successfully executing critical rare state-action pair associated with sparse reward. The indicator function $\mathds{1}(\cdot)$ evaluates to $1$ when condition $f_{C}(s_{t},a_{t})={True}$ satisfied, indicating the successful execution of action $a_{t}$ in state $s_{t}$. The stage reward term is modulated by a time-dependent fading function $f{\phi}(t)$. These extrinsic rewards obtained from the environment are shared by human and AI agents.
@@ 4.1.2 Intrinsic Reward Design
a) AI self-motivated intrinsic reward.
To encourage the AI agent to adopt a diverse policy that pays more attention to rare state-action pairs, we design self-motivated intrinsic reward as
$\begin{split}r_{t}^{\mathrm{A}}&=\lambda^{\mathrm{A}}\cdot\mathbb{E}_{\pi}\left[-\log\left(\pi(a_{t}^{\mathrm{A}}\big{|}o_{t}^{\mathrm{A}})\right)\right],\end{split}$ (3)
@@
Since capturing the sparse rewards in HAIC relies on thorough explorations of the critical rare state-action pairs associated with sparse rewards, our formulation given in eq. (3) differs from classical entropy-based intrinsic rewards, typically expressed by $\hat{r}_{t}^{\mathrm{A}}=\lambda^{\mathrm{A}}\cdot\mathbb{E}_{\pi}[-{\pi}(a_{t}^{\mathrm{A}}|o_{t}^{\mathrm{A}})\log{\pi}(a_{t}^{\mathrm{A}}|o_{t}^{\mathrm{A}})]$. By eliminating the probability term, ${\pi}(a_{t}^{\mathrm{A}}|o_{t}^{\mathrm{A}})$, in the classical entropy formulation, our logarithmic formulation provides stronger motivations for exploring rare state-action pairs. This design principle also extends to our human-motivated intrinsic reward component detailed in the following.
b) Human-motivated intrinsic reward.
Our human-motivated intrinsic reward is defined as
$r_{t}^{\mathrm{H}}=\lambda^{\mathrm{H}}\cdot\mathbb{E}_{\pi}\left[\left|\log\left(\frac{\pi\left(a_{t}^{\mathrm{A}}\big{|}o_{t}^{\mathrm{A}}\right)}{{\pi}\left(\tilde{a}_{t}^{\mathrm{A}}\big{|}\tilde{o}_{t}^{\mathrm{A}}\left(a_{t}^{\mathrm{H}},o_{t}^{\mathrm{A}}\right)\right)}\right)\right|\right],$ (4)
@@
We extend the logarithmic formulation from the AI’s self-motivated intrinsic reward by incorporating an absolute value function ||$\cdot$|| ⋅ |. This modification ensures equal treatment of both overestimated (where the AI agent shows higher action probability in actual scenarios) and underestimated (where the AI agent shows higher action probability in counterfactual scenarios) behavioral impacts, preventing biased reward signals that could impair exploration. Through symmetric treatment of these behavioral deviations, the reward mechanism effectively amplifies significant discrepancies between actual and counterfactual behaviors, enabling the AI agent to develop policies that robustly adapt to and complement human actions.
These two intrinsic rewards given in eqs. (3) and (4) encourage the AI agent to explore actions that can increase the likelihood of encountering the critical rare state-action pairs associated with sparse rewards. This strategic exploration boosts the frequency of targeted behaviors, helping the AI agent adapt to human behavior effectively in HAIC.
@@ 4.1.3 Context-aware Weights Design
$\begin{split}\bar{R}_{n}^{\mathcal{E}}=\frac{1}{T}\sum_{t=nT}^{(n+1)T}r_{t}^{\mathcal{E}},~{}~{}~{}\bar{R}_{n}^{\mathrm{A}}=\frac{1}{T}\sum_{t=nT}^{(n+1)T}r_{t}^{\mathrm{A}},~{}~{}~{}\bar{R}_{n}^{\mathrm{H}}&=\frac{1}{T}\sum_{t=nT}^{(n+1)T}r_{t}^{\mathrm{H}},\end{split}$ (7)
@@
$r_{t}^{\mathcal{E}}=r_{t}^{\mathrm{E}_{\mathrm{S}}}+r_{t}^{\mathrm{E}_{\mathrm{G}}}$ (8)
@@ 4.2 Training Algorithm
Our BCR-DRL agent is adapted from PPO Schulman et al. (2017). Its architecture consists of actor and critic neural networks with parameters $\theta$ and $\varphi$, respectively. The actor generates actions $a_{t}^{\mathrm{A}}$ using a stochastic policy $\pi(a_{t}^{\mathrm{A}}|o_{t}^{\mathrm{A}};\theta)$, which represents the probability density of $a_{t}^{\mathrm{A}}$ given the current state $o_{t}^{\mathrm{A}}$. The critic estimates the state-value function $V(o_{t}^{\mathrm{A}};\varphi)$ based on the actor’s policy. Please refer to Algorithm 1 for the step-by-step training process of BCR-DRL for HAIC.
@@ 5 Experiments
As introduced in Section 3, we provide experiments in the overcooked environment Carroll et al. (2019), which presents cooking tasks to be accomplished by coordinated actions of human and AI agents within limited timesteps. Specifically, a human agent and an AI agent work together to prepare as many onion soups as possible within a limited number of timesteps. A shared sparse reward ($r_{t}^{\mathrm{E}}$ given in eq. (2)) of $20$ points will be granted for both agents upon each successfully served onion soup, which requires completing a sequence of actions, including picking up onions from specified locations, placing three onions into a pot, picking up the soup with a plate after cooking it for $20$ timesteps, and serving the soup to a designated area.
We validate the effectiveness of our BCR-DRL in three of the Overcooked layouts. focusing on different aspects of HAIC. In our experiments, we focus on model-based HAIC, using human behavioral models trained using behavior cloning algorithms and data is sourced from Carroll (2024); Carroll et al. (2019). Our implementation is based on the Gym-compatible Overcooked environment given in using Tensorflow Carroll (2024).
Figure 2: Average sparse rewards of the Cramped Room layout, obtained by the training process in (i) and the testing results in (ii).
@@
Figure 3: Impacts of the proposed methods in the Cramped Room layout. The intrinsic reward scheme is given in (a), whilst the context-aware weighting mechanism is given in (b). Specifically, the Rewards in (a) correspond to three algorithms: PPOBC (i), Causal (ii), and BCR-DRL (iii); whilst the Weights (b) compose the three context-aware weights of the BCR-DRL given in eq. (5), representing the environment feedback weight ${\kappa}_{n}^{\mathcal{E}}$ (i), the AI behavior diversity weight ${\kappa}_{n}^{\mathrm{A}}$ (ii), and the human coordinator behavior weight ${\kappa}_{n}^{\mathrm{H}}$ (iii), respectively
5.1 Benchmark Algorithms
We evaluate our BCR-DRL algorithm against two benchmark algorithms. All three methods utilize the standard proximal policy optimization (PPO) framework Schulman et al. (2017); Chrysovergis (2024), with a pre-trained human agent.
@@ 5.1 Benchmark Algorithms
For fair comparisons, all three algorithms share the same neural network architecture and hyper-parameter settings.
5.2 Results and Analysis
For each experimental layout, we provide statistical experimental training and testing results, based on $5$ independent training runs and $2000$ independent testing episodes, respectively.
5.2.1 Cramped Room Layout
In the Cramped Room layout, the human and AI agents operate in a shared space, necessitating precise movement planning policies to avoid collisions. We analyze our method’s performance from the following two perspectives given in Figs. 3 and 3, respectively.
Sparse reward performance comparison.
As shown in Fig. 3, our BCR-DRL consistently outperforms benchmark methods, and achieves approximately $20\%$ higher average sparse rewards in both training and testing phases. While the PPOBC benchmark provides a solid foundation through its focus on the extrinsic rewards, and the Causal benchmark offers valuable insights through counterfactual actions of the ego AI agent, our BCR-DRL builds upon these approaches by designing a more effective and efficient solution for HAIC in this shared-space HAIC scenario.
Effectiveness of BCR design.
To understand the attributes behind BCR-DRL’s performance, we analyze experimental results regarding its key components in Fig. 3, regarding both intrinsic rewards in Fig.3(a) and context-aware weights in Fig. 3(b).
Fig.3(a) shows the contribution of different reward components across all three algorithms: PPOBC, Causal, and BCD-DRL. The notably higher intrinsic rewards in BCR-DRL compared to the Causal benchmark indicate a more comprehensive exploration of human behavior patterns, demonstrating the effectiveness of our dual intrinsic design.
Fig.3(b) illustrates the temporal evolution of context-aware weights during training, revealing two distinct patterns. In the initial phase, higher weights are allocated to human-motivated intrinsic rewards alongside extrinsic rewards, reflecting the critical importance of interpreting human partner behaviors for collision avoidance in the environment. As training progresses, interestingly, the weights of AI self-motivated intrinsic rewards become dominant, facilitating a thorough exploration of critical but rare state-action pairs, which promotes more targeted exploration in sparse reward scenarios. Concurrently, the weights of extrinsic rewards diminish, indicating the system’s emphasis on exploration driven by intrinsic rewards. These observations in Fig.3(b) demonstrate that BCR-DRL’s adaptive weighting mechanism adaptively balances exploration and exploitation, ultimately enabling efficient Human-AI Collaboration (HAIC).
Figure 4: Average sparse rewards of the Asymmetric Advantages layout, obtained by the training process in (i) and the testing results in (ii).
Figure 5: Impacts of the proposed methods in the Asymmetric Advantages layout. The intrinsic reward scheme is given in (a), whilst the context-aware weighting mechanism is given in (b).
5.2.2 Asymmetric Advantages Layout
The Asymmetric Advantages layout presents a unique coordination challenge where human and AI agents operate from separate areas and have varying access to cooking resources. Unlike the Cramped Room layout where human and AI agents share the same space and cooking resources, this Asymmetric Advantages layout creates natural accessibility disparities. For example, agents would have different distances for serving the soup to the destinations in their located area from the same pot. This layout emphasizes the HAIC from the perspective of leveraging agents’ asymmetric advantages.
Sparse reward performance comparison.
Fig. 5 shows that our BCR-DRL achieves consistently higher sparse rewards compared to both benchmarks, during both training and testing phases. The algorithm’s effectiveness in this layout is particularly evident in its convergence speed, reaching optimal performance approximately $67\%$ faster than baseline methods.
Interestingly, for the testing results, Fig. 5 shows that the sparse rewards achieved by the Causal benchmark are significantly lower compared with the other two algorithms. This observation underscores the robustness side of our context-aware weighting mechanism in HAIC, particularly when intrinsic rewards are integrated to enhance DRL exploration.
Effectiveness of BCR design.
The experimental results in Fig. 5 provide deeper insights into BCR-DRL’s superior performance in asymmetric resource scenarios. The evolution of intrinsic rewards and context-aware weights reveals two distinct characters of BCR in Figs. 5(a) and 5(b), respectively.
Fig. 5(a) demonstrates reward component trends similar to those observed in the Cramped Room layout. In addition to a more significant value of the intrinsic rewards compared with that in the Causal benchmark, we can also observe a notable inflection point in the training results of Fig. 5(a)(iii), caused by context-aware weights effectively truncate the influence of intrinsic rewards when exploitation needs be to focused.
Fig. 5(b) reveals that context-aware weights for AI self-motivated intrinsic rewards increase significantly during the initial training phase, contrasting with the Cramped Room layout where other rewards dominate early training. This distinction arises from two factors in the Asymmetric Advantages layout: the initial scarcity of human behavior data and the spatial separation of agents with differential access to cooking resources. Unlike the Cramped Room scenario, where agents operate in close proximity, the physical separation in this layout reduces the immediate impact of human actions on the AI agent’s performance. Consequently, prioritizing the AI agent’s role adaptation through AI self-motivated exploration proves more effective when human behavioral data is limited. As training progresses and human behavior data accumulates, the weights of human-motivated intrinsic rewards gradually increase, enhancing coordination in HAIC scenarios by leveraging asymmetric information sources for HAIC optimization.
5.2.3 Coordination Ring Layout
The Coordination Ring layout introduces a unique HAIC challenge where agents navigate around a central obstacle in a small room. This design creates a scenario where agents must carefully plan their movements in a confined space.
Sparse reward performance comparison.
@@ 5.2.3 Coordination Ring Layout
As shown in Fig. 7, while BCR-DRL still outperforms the benchmarks, the performance improvement is less pronounced compared to the previous two layouts. This can be attributed to the confined nature of the Coordination Ring layout, which naturally constrains the state space and enables the PPOBC benchmark to achieve relatively comprehensive exploration even without sophisticated reward mechanisms. An interesting observation is that the Causal benchmark exhibits higher variance compared to both BCR-DRL and PPOBC, suggesting that purely partner behavior-motivated intrinsic rewards may be less stable in environments with a confined space that is easy to explore comprehensively.
Effectiveness of BCR design.
Fig. 7(a) reveals distinct patterns of the reward components specific to the Coordination Ring layout. The Causal benchmark shows significantly larger intrinsic rewards with high variance, explaining its unstable sparse rewards in this constrained environment in Fig. 7. Unlike the previous two layouts, the human-motivated intrinsic reward of BCR-DRL here maintains a relatively consistent level throughout training without a clear inflection point, suggesting that the confined space naturally encourages more intensive human-AI interactions from the early stages of learning.
In Fig. 7(b), the evolution of context-aware weights also exhibits a distinct pattern in this layout. We can observe that all three weights are comparably stable compared with those in the other two layouts, except for those in the very early training stage, we observe slightly lower environmental feedback weights coupled with higher AI behavior diversity weights, indicating that BCR-DRL initially prioritizes developing diverse movement strategies compared with explore through the environment.
Figure 6: Average sparse rewards of the Coordination Ring layout, obtained by the training process in (i) and the testing results in (ii).
Figure 7: Impacts of the proposed methods in the Coordination Ring layout. The intrinsic reward scheme is given in (a), whilst the context-aware weighting mechanism is given in (b).
These observations in the Coordination Ring layout suggest that while traditional PPOBC demonstrates competitive performance in confined environments that are relatively straightforward to explore comprehensively, BCR-DRL offers a more robust and generalizable approach, capable of adapting to the diverse challenges inherent in HAIC scenarios.
6 Conclusion
To enhance HAIC, we introduced a BCR-DRL to facilitate HAIC. Supplementing traditional extrinsic rewards, our BCR incorporated an innovative dual intrinsic rewarding scheme to facilitate comprehensive exploration and a novel context-aware weighting mechanism to optimize exploration and exploitation. Extensive experimental results across three layouts of the Overcooked environment demonstrated that BCR-DRL can increase rewards by approximately $20\%$ and reduce training time by approximately $67\%$. Testing experiments underscored the algorithm’s robustness and generalization abilities across different coordination aspects of human and AI.

### v5

@@ 4 The Proposed BCR-DRL Algorithm
In this section, we first present the design specifics of our BCR, followed by the training algorithm of BCR-DRL for HAIC.
@@ 4.1 Design Specifics of BCR
The dual intrinsic rewards encourage the exploration of critical rare state-action pairs that are associated with sparse rewards, capturing distinctive behavioral patterns from both the AI agent and its human coordinator. The context-aware weights adaptively adjust each reward component’s significance epoch by epoch, maintaining synchronization with the BCR-DRL policy updates. We note that the use of distinguishable superscripts ($\mathrm{E}$ and $\mathcal{E}$) for the extrinsic reward and its context-aware weight, is intentional to emphasize their distinct design rationale, which will be elaborated in eq. (11).
@@ 4.1.1 Extrinsic Reward
The extrinsic reward, $r_{t}^{\mathrm{E}}$, comprises two components: the target sparse reward, $r_{t}^{\mathrm{E}_{\mathrm{S}}}$, and a stage reward [^1] Stage rewards are defined as the intermediate rewards used to guide the exploration and exploitation of key preliminary actions that lead to the target sparse rewards [24, 25]. We follow the methods of stage rewards designing given in the supplementary of [1], thereby saving the space in this paper for discussing its effectiveness., $r_{t}^{\mathrm{E}_{\mathrm{G}}}$. This extrinsic reward is given by
$r_{t}^{\mathrm{E}}\!\!=\!r_{t}^{\mathrm{E}_{\mathrm{S}}}+r_{t}^{\mathrm{E}_{\mathrm{G}}}\!\cdot\!f_{\phi}(t)\!\!=\!\!\lambda^{\mathrm{E}_{\mathrm{S}}}\!\cdot\!\mathds{1}\left\{f_{C}(s_{t},a_{t})\!\!=\!{True}\right\}+r_{t}^{\mathrm{E}_{\mathrm{G}}}\!\cdot\!f_{\phi}(t),$ (5)
where $\lambda^{\mathrm{E}_{\mathrm{S}}}$ is a constant coefficient representing the magnitude of successfully executing critical rare state-action pair associated with sparse reward. The indicator function $\mathds{1}(\cdot)$ evaluates to 111 when condition $f_{C}(s_{t},a_{t})={True}$ satisfied, indicating the successful execution of action $a_{t}$ in state $s_{t}$. The stage reward term is modulated by a time-dependent fading function $f_{\phi}(t)$. These extrinsic rewards obtained from the environment are shared by human and AI agents.
@@ 4.1.2 Intrinsic Reward Design
AI self-motivated intrinsic reward. To encourage the AI agent to adopt a diverse policy that pays more attention to rare state-action pairs, we design the AI self-motivated intrinsic reward following Theorem 1 as
$\begin{split}r_{t}^{\mathrm{A}}&=\lambda^{\mathrm{A}}\cdot\mathbb{E}_{\pi}\left[\left|\log\left(\pi(a_{t}^{\mathrm{A}}\mid o_{t}^{\mathrm{A}})\right)\right|\right],\end{split}$ (6)
@@
Human-motivated intrinsic reward. We design the human-motivated intrinsic reward based on Theorem 1, which is given by
$r_{t}^{\mathrm{H}}=\lambda^{\mathrm{H}}\cdot\mathbb{E}_{\pi}\left[\left|\log\left(\frac{\pi\left(a_{t}^{\mathrm{A}}\mid o_{t}^{\mathrm{A}}\right)}{{\pi}\left(\tilde{a}_{t}^{\mathrm{A}}\mid\tilde{o}_{t}^{\mathrm{A}}\left(a_{t}^{\mathrm{H}},o_{t}^{\mathrm{A}}\right)\right)}\right)\right|\right],$ (7)
@@
Two intrinsic rewards given in eqs. (6) and (7) encourage the AI agent to explore actions that can increase the likelihood of encountering the critical rare state-action pairs associated with sparse rewards. This strategic exploration boosts the frequency of targeted behaviors, helping the AI agent adapt to human behavior effectively in HAIC.
@@ 4.1.3 Context-aware Weights Design
$\begin{split}\bar{R}_{n}^{\mathcal{E}}=\frac{1}{T}\hskip-6.0pt\sum_{t=nT}^{(n+1)T}\hskip-7.0ptr_{t}^{\mathcal{E}},~~~\bar{R}_{n}^{\mathrm{A}}=\frac{1}{T}\hskip-6.0pt\sum_{t=nT}^{(n+1)T}\hskip-7.0ptr_{t}^{\mathrm{A}},~~~\bar{R}_{n}^{\mathrm{H}}&=\frac{1}{T}\hskip-6.0pt\sum_{t=nT}^{(n+1)T}\hskip-7.0ptr_{t}^{\mathrm{H}},\end{split}$ (10)
@@
$r_{t}^{\mathcal{E}}=r_{t}^{\mathrm{E}_{\mathrm{S}}}+r_{t}^{\mathrm{E}_{\mathrm{G}}},$ (11)
@@ 4.2 Training Algorithm
Our BCR-DRL agent is adapted from proximal policy optimization (PPO) [26]. Its architecture consists of actor and critic neural networks with parameters $\theta$ and $\varphi$, respectively. The actor generates actions $a_{t}^{\mathrm{A}}$ using a stochastic policy $\pi(a_{t}^{\mathrm{A}}|o_{t}^{\mathrm{A}};\theta)$, which represents the probability density of $a_{t}^{\mathrm{A}}$ given the current state $o_{t}^{\mathrm{A}}$. The critic estimates the state-value function $V(o_{t}^{\mathrm{A}};\varphi)$ based on the actor’s policy. Please refer to Algorithm 1 for the step-by-step training process of BCR-DRL for HAIC.
(a) Cramped Room.
(b) Asymmetric Advantages.
(c) Coordination Ring.
Figure 2: Average sparse rewards in different layouts of the Overcooked environment, obtained by training in (i) and testing in (ii).
@@ 5 Experiments
Following the original Overcooked environment setup [1], our experimental framework consists of one human model and one RL-based AI agent, making our approach effectively a single-agent reinforcement learning problem where we train a single-agent PPO to coordinate with the human model. We note that while the training process uses a combination of multiple reward components (extrinsic and intrinsic rewards), we report only the sparse environmental rewards during both training and evaluation phases, as they serve as the primary metric for task completion performance. The detailed training curves for individual reward components are provided in Fig. 4.
We evaluate BCR-DRL algorithm against two benchmark algorithms: 1) PPOBC benchmark: This benchmark uses only the extrinsic reward, without incorporating intrinsic rewards or context-aware weights adjustments [1]. 2) Causal benchmark: This benchmark includes causal influence rewards in addition to the extrinsic rewards used in PPOBC benchmark. We include this benchmark to analyze the effectiveness of intrinsic rewards that solely focus on social influence as intrinsic motivation proposed in [10], without incorporating context-aware weights. All algorithms utilize the standard PPO framework [26, 27], and are evaluated with corresponding human models tailored for individual scenarios. For fair comparisons, all algorithms share the same neural network architecture and hyperparameter settings in the same scenario.
(a) Cramped Room.
(b) Asymmetric Advantages.
(c) Coordination Ring.
Figure 3: Ablation studies comparing our BCR-DRL full model against two variants, BCR-NoIntrinsic and BCR-NoWeights. The comparison between BCR-DRL and BCR-NoIntrinsic demonstrates the effectiveness of intrinsic rewards in improving performance. The comparison between BCR-DRL and BCR-NoWeights shows the stabilizing effect of our context-aware weighting mechanism, which mitigates potential instabilities from intrinsic rewards by preventing them from dominating later-stage learning.
(a) Cramped Room.
(b) Asymmetric Advantages.
(c) Coordination Ring.
@@
Figure 4: Performance analysis of individual components in BCR. Impacts of the proposed methods in the Cramped Room, Asymmetric Advantages, and Coordination Ring layouts, respectively. The intrinsic rewards are given in (i)-(iii) of each figure, whilst the context-aware weights are given in (iv)-(vi). Specifically, the Rewards correspond to three algorithms: PPOBC (i), Causal (ii), and BCR-DRL (iii); whilst the Weights compose the three context-aware weights of the BCR-DRL given in eq. (8), representing the environment feedback weight ${\kappa}_{n}^{\mathcal{E}}$ (iv), the AI behavior diversity weight ${\kappa}_{n}^{\mathrm{A}}$ (v), and the human coordinator behavior weight ${\kappa}_{n}^{\mathrm{H}}$ (vi), respectively.
5.1 Performance on Overcooked Environment
The overcooked environment [1] presents cooking tasks to be accomplished by coordinated actions of human and AI agents within limited timesteps. Specifically, a human agent and an AI agent work together to prepare as many onion soups as possible within a limited number of timesteps. A shared sparse reward ($r_{t}^{\mathrm{E}}$ given in eq. (5)) of 202020 points will be granted for both agents upon each successfully served onion soup, which requires completing a sequence of actions, including picking up onions from specified locations, placing three onions into a pot, picking up the soup with a plate after cooking it for 202020 timesteps, and serving the soup to a designated area.
We evaluate our BCR-DRL approach across three distinct Overcooked layouts, each highlighting different aspects of HAIC. Our experiments employ model-based HAIC using human models trained through behavior cloning algorithms. Each layout utilizes different sets of real human interaction data sourced from [22, 1], with separate human models used during training and testing phases to investigate robustness. Our implementation leverages the Gym-compatible Overcooked environment with Tensorflow [22].
To ensure statistical reliability and demonstrate the stability of our approach, including the context-aware weighting mechanism, all experimental results are based on comprehensive statistical analysis across multiple independent runs. Specifically, for each experimental layout, we provide statistical experimental training and testing results, based on 555 independent training runs and 200020002000 independent testing episodes, respectively. This rigorous experimental protocol ensures that our reported performance improvements and the stability of context-aware weights are statistically significant and reproducible across different training conditions.
5.1.1 Sparse Reward Performance Comparison
Fig. 2 shows the comparative performance of sparse rewards between our BCR-DRL and the benchmark algorithms.
@@ 5.1.1 Sparse Reward Performance Comparison
Coordination Ring. As shown in Fig. 2(c), while BCR-DRL still outperforms the benchmarks, the performance improvement is less pronounced compared to the previous two layouts. This can be attributed to the confined nature of the Coordination Ring, which naturally constrains the state space and enables the PPOBC benchmark to achieve relatively comprehensive exploration even without sophisticated reward mechanisms. An interesting observation is that the Causal benchmark exhibits higher variance compared to both BCR-DRL and PPOBC, suggesting that purely partner behavior-motivated intrinsic rewards may be less stable in environments with a confined space that is easy to explore comprehensively. These observations in the Coordination Ring layout suggest that while traditional PPOBC demonstrates competitive performance in confined environments that are relatively straightforward to explore comprehensively, BCR-DRL offers a more robust and generalizable approach, capable of adapting to the diverse challenges inherent in HAIC scenarios.
5.1.2 Ablation Studies
Our ablation study examines BCR’s key components across three Overcooked layouts by comparing our full model against: BCR-NoIntrinsic (without intrinsic rewards) and BCR-NoWeights (with intrinsic rewards but without context-aware weights). Fig. 3 demonstrates that component impact varies according to the coordination requirements of each layout. In the Cramped Room (Fig. 3(a)), BCR-NoWeights initially matches the full model but becomes unstable after 100 epochs, confirming our design intuition from Section 4.1.3 that unweighted intrinsic rewards become detrimental during exploitation phases. For Asymmetric Advantages (Fig. 3(b)), BCR-NoWeights performs adequately but underperforms during later training stages when balanced reward emphasis becomes critical. In the Coordination Ring (Fig. 3(c)), all variants achieve similar performance, indicating that confined spaces require less sophisticated exploration. Note that BCR-NoIntrinsic is equivalent to the PPOBC benchmark (see Section 5.1.1 for detailed analysis). These ablation studies validate that dual intrinsic rewards enhance critical exploration in the early training stages, while context-aware weights improve stability during exploration-exploitation transitions, with benefits most significant in scenarios requiring comprehensive exploration and extensive agent coordination.
5.1.3 Effectiveness Analysis on BCR Design
As shown in Fig. 4, we analyze the effectiveness of BCR’s key components through their training curves. We recall that the definitions of all components are given in Section 4.1.
Cramped Room. In Fig. 4(a)(ii)-(iii), BCR-DRL consistently outperforms the Causal benchmark in intrinsic rewards, demonstrating enhanced exploration of human behavior patterns. The component weight evolution in Figs. 4(a)(iv)-(vi) reveals a systematic adaptation process. During early training, environment feedback (iv) and behavior diversity weights (v) dominate as the AI effectively leverages direct environmental information in this shared space. As training progresses, human coordinator behavior weights (vi) gradually increase, reflecting the AI’s focus shifting with the increase in human behavioral data. This adaptive weighting illustrates how the algorithm efficiently transitions from environmental exploration to HAIC in a shared space.
Asymmetric Advantages. Fig. 4(b)(iii) shows human-motivated intrinsic rewards with a distinctive inflection point, reflecting the context-aware weighting mechanism’s ability to modulate reward signals during the exploration-exploitation transition. Unlike in Cramped Room, the human coordinator behavior weights (vi) reach significant values earlier in the training process. This stems from the environment’s spatial separation and asymmetric resource distribution, which make human behaviors informative for policy learning. The adaptive process reflects how BCR-DRL improves AI’s adaptation in environments with spatial differences and asymmetric access to critical resources.
Coordination Ring. Fig. 4(c) shows distinct patterns, with the Causal benchmark exhibiting significantly larger intrinsic rewards with high variance, explaining its unstable performance in Fig. 2(c). BCR-DRL’s human-motivated intrinsic reward maintains relatively consistent levels throughout training, suggesting this confined space is naturally easier to explore, limiting BCR’s exploration advantages. Context-aware weights demonstrate greater stability compared to other layouts, with slightly lower environmental feedback weights and higher AI behavior diversity weights initially, indicating BCR-DRL prioritizes diverse movement strategies over environment exploration in early training.
5.2 Hyperparameter Selection and Sensitivity Methodology
We present a systematic analysis of BCR-DRL’s critical hyperparameters: intrinsic reward truncation epoch ($N_{\mathrm{th}}$), intrinsic reward coefficients ($\lambda^{\mathrm{A}},\lambda^{\mathrm{H}}$), and softmax temperature ($\lambda^{\mathrm{R}}$). Complete configurations are provided in our code repository at https://github.com/hxheart/BCR-DRL.
Intrinsic Reward Truncation Epoch ($N_{\mathrm{th}}$). This parameter in eq. (8) governs the exploration-exploitation transition. The optimal value correlates with sparse reward improvement rates: rapidly increasing rewards indicate insufficient exploration, requiring later truncation, while plateaued rewards suggest earlier transition to exploitation. For Cramped Room, we set $N_{\mathrm{th}}=100$ based on the training curve inflection point (Fig. 2(a)), as Fig. 3(a) shows that $N_{\mathrm{th}}=150$ leads to performance degradation from over-exploration. Following similar analysis, we set $N_{\mathrm{th}}=150$ for Asymmetric Advantages and $N_{\mathrm{th}}=300$ for Coordination Ring. This parameter is sensitive to environment-specific reward dynamics, as evidenced by the wide variation in optimal values across environments (100-300 epochs), requiring calibration rather than using a universal setting.
Intrinsic Reward Coefficients ($\lambda^{\mathrm{A}}$, $\lambda^{\mathrm{H}}$). These coefficients in eqs. (6) and (7) calibrate AI self-motivated and human-motivated intrinsic rewards. Optimal configuration maximizes exploration while maintaining effective exploitation. We maintain $\lambda^{\mathrm{A}}>\lambda^{\mathrm{H}}$ to compensate for lower human action frequency compared to AI actions, balancing both agents’ behavioral influence. For Cramped Room and Asymmetric Advantages, empirical testing established $\lambda^{\mathrm{A}}=1$ and $\lambda^{\mathrm{H}}=0.02$. The spatially confined Coordination Ring, which facilitates inherent exploration even in traditional PPO, benefits from reduced coefficients: $\lambda^{\mathrm{A}}=0.2$ and $\lambda^{\mathrm{H}}=0.01$. The coefficients show sensitivity to environment characteristics, particularly spatial constraints, though the relative ratio $\lambda^{\mathrm{A}}>\lambda^{\mathrm{H}}$ remains consistently effective across different scenarios.
Softmax Temperature ($\lambda^{\mathrm{R}}$): This parameter in eq. (9) ensures smooth adaptation among all three reward components in eq. (4). We determined $\lambda^{\mathrm{R}}=3$ to allow each component to maintain proportional influence throughout training. This parameter demonstrates low sensitivity to environment variations, with the uniform value across all tested environments indicating robust performance without environment-specific tuning.
5.3 Generalization Study
To evaluate the generalization capability of BCR-DRL beyond the Overcooked domain and the behavior-cloned human models used in previous experiments, we introduce a new environment, Exploration, which features a distinct task setting and a different human-agent behavior model.
As shown in Fig. 5(a), this environment consists of a grid populated with unexplored cells (blank), explored cells (green), and obstacles (black). A human agent (blue) and an AI agent (red) jointly explore all accessible cells in the environment. Agents can choose from five discrete actions: move up, down, left, right, or stay. The environment is episodic, with each episode spanning 400 timesteps, potentially containing multiple rounds of exploration. Unlike prior experiments that relied on human-cloned policies learned from real gameplay data, we employ a synthetic human model with a stochastic policy and reduced temporal frequency (acting every 10 timesteps) based on real human reaction time. This introduces significant behavioral uncertainty and temporal asynchrony, creating a coordination scenario that differs from the Overcooked setups. The AI agent, by contrast, acts at every timestep and must adaptively coordinate its exploration strategy to complement the human’s delayed and unpredictable actions. Please also refer to Section 3.2 for the analysis of HAIC aspects of this Exploration environment. The reward structure incentivizes efficient and cooperative behavior: agents receive +2 for discovering a new cell, -0.5 for revisiting an already explored cell, and -1 for invalid actions (including boundary violations, collisions with obstacles or the other agent, or redundant actions with no position change). When all accessible cells are explored, both agents receive a +20 sparse reward, and the environment resets while retaining the obstacle layout.
This setup challenges the AI agent to anticipate human movements, avoid redundancy, and dynamically adapt its behavior under conditions of uncertainty and limited observability. It thereby provides a controlled but difficult testbed for evaluating whether BCR-DRL can generalize to novel coordination dynamics and behavior patterns that diverge from its training conditions. Experimental results in Fig. 5(b) show that BCR-DRL consistently outperforms the benchmark approaches (PPOHM and Causal), achieving higher cumulative rewards and sample efficiency. These findings validate BCR-DRL’s effectiveness in generalizing across environments and partner behavior models, reinforcing its potential as a robust solution for real-world HAIC applications.
(a) Exploration env.
(b) Averaged sparse rewards.
Figure 5: Performance comparison based on BCR-DRL and the benchmark algorithms on the Exploration environment.
6 Conclusion, Limitations, and Future Work
[…]